# Towards Context-Aware Domain Generalization: Representing Environments with Permutation-Invariant Networks

### 15 Dec 2023 v1

@@ 3.2 Learning Permutation-Invariant Representations
Notably, the methods above impose no probabilistic structure on the set-representations, since the latter are mainly used as deterministic features for downstream tasks. In contrast, probabilistic models attempt to learn a conditional or a marginal distribution over the set-representations. In the Bayesian literature, sets represent finitely exchangeable sequences, embodying the core probabilistic structure of most Bayesian models [44]. In particular, hierarchical or multi-level Bayesian models are used to model the dependencies in nested data, where observations are organized into clusters or levels [64, 21], mirroring the notions of domain or environment in DG. Indeed, hierarchical Bayesian models have been successfully applied in many areas of science, but are typically constrained to linear or generalized linear models (GLMs), prioritizing interpretability over predictive performance.
From a somewhat different Bayesian perspective, IID data represents the starting point for learning invariant summary statistics for parameter estimation or model comparison. The pioneering work on Neural Statisticians [14] tackles the task from a variational perspective, learning global set-representations as part of a generative model. Neural processes [19, 30] comprise a related family of set-based models for prediction and uncertainty quantification in supervised learning tasks, rooted in the spirit of Gaussian processes [48]. Finally, [31] explore a permutation-invariant variational autoencoder (SetVAE) with multiple latent variables trained with a modified ELBO objective. The goal of this and follow-up models [71] is accurate generative performance, less so the learning of compact representations. The theoretical and empirical implications of learning random set-representations falls outside the scope of our current work, but offers a potentially fruitful avenue for future research.
Crucially, none of the above methods pursues the concrete goal of our work, which is improving domain generalization performance through set-based environment representations. For the purpose of achieving this goal, we employ variants of the DeepSet [69] and SetTransformer architectures for our backbone set-encoder throughout all experiments. Interestingly, we observe that our approach is widely robust to the choice of set-encoder architectures in the problems considered.
3.3 OOD Detection and Selective Classification
Detecting unusual inputs that deviate from the examples in the training set has been a long-standing problem of conceptual complexity in machine and statistical learning [3, 67, 52, 23, 66]. Flagging out-of-distribution (OOD) instances involves identifying uncommon data points that might compromise the reliability of machine learning systems [67]. OOD detection is closely related to inference with a reject option (also termed selective classification) [20, 16], which allows classifiers to refrain from making a prediction under ambiguous or novel conditions [25]. The reject option has been extensively studied in statistical and machine learning [24, 18, 22, 62], with early work dating back to the 1950s[11, 10, 24].
More recently, [41] explored the utility of selective classification in DG settings. They investigated various post-hoc scores to define a “competence region” in feature space where a classifier is deemed competent. Post-hoc methods rely on various aspects of pre-trained model outputs, such as the softmax outputs (e.g., [28]), logit outputs (e.g., [26, 36]), or intermediate features (e.g., [2, 53, 59]). In this work, we consider a post-hoc score based on the $k$-nearest neighbors to the training set in feature space similar to [53], which is applicable to both classification and regression settings. Unlike the approach taken in [41], where the focus lies on individual input features, we consider the set summary provided by the set-encoder. Thus, we can identify novel environments even when singleton inputs lack sufficient information.
4 Experiments
Figure 2: Experiment 1. Left: Toy dataset that conforms to our theoretical criteria. A detailed description can be found in Section C.1. Without environmental information, the marked input at $x=2.5$ could belong to either one of the domains numbered 1, 2, or 3. Marginal distributions of the five environments are shown on top. Right: Comparison of environment classification accuracy for different set sizes.
In the following, we explore various aspects of our approach across three different dimensions. First, we show on two datasets that our model achieves improved performance in the ID as well as the OOD setting compared to a baseline model. Second, we show how novel environments can be detected to select between the most predictive (in the ID setting) and the most robust (in the OOD setting) model. We also show that novel environment detection can be utilized to avoid failure cases. Third, we demonstrate that the necessary criteria (see Section 2.5) can be validated empirically, identifying cases where no benefits of our method can be expected. Experimental details can be found in the Appendix.
4.1 Evaluation Approach
Here, $\mathcal{L}(f^{\mathbf{Y}\mid\mathbf{X},\mathbf{S}^{(n)}})$ denotes a performance measure for our model and similarly for the baseline model $f^{\mathbf{Y}\mid\mathbf{X}}$. $\mathcal{R}_{\text{I}}>0$ signifies an advantage attained by our approach and therefore the fulfillment of Criterion 2.1. In the regression setting, we consider the negative L2-Loss as the performance measure.
To validate Criterion 2.2, we train a contextual environment model (referred to as $f^{E\mid\mathbf{X},\mathbf{S}^{(n)}}$) utilizing both the set input $\mathbf{S}^{(n)}$ and the target input $\mathbf{X}$ to predict the environment label $E$. Additionally, we train a baseline environment model (denoted as $f^{E\mid\mathbf{X}}$) aimed at predicting $E$ solely from $\mathbf{X}$. We then compute the relative improvement $\mathcal{R}_{\text{II}}$ [^3] The exact definition for $\mathcal{R}_{\text{II}}$ and $\mathcal{R}_{\text{III}}$ can be found in Appendix B of the contextual environment predictor relative to the baseline environment predictor. $\mathcal{R}_{\text{II}}>0$ indicates that Criterion 2.2 is satisfied. In our experiments, we choose the set size $n$ such that we achieve approximately 100% accuracy for our contextual environment predictor $f^{E\mid\mathbf{X},\mathbf{S}^{(n)}}$ on ID data.
4.2 Experiment 1: Toy Example
@@
To set the stage, we consider a dataset that is inspired by [63]. The dataset includes data from five environments, defined by five 2D Gaussian distributions differing only in their locations (i.e., mean vectors). The data-generating process is thus:
$j$ $\sim\text{Categorical}(p_{1},\dots,p_{J})$ (5)
$\mathbf{X}^{(ij)}$ $\sim\mathcal{N}(\boldsymbol{\mu}^{(j)},\mathbf{\Sigma})\quad\text{for}\quad i=1,\dots,I_{j}$ (6)
This is the classical setting of Simpsons’ paradox, for which a naive linear model fit oblivious to the hierarchical data-generating process will yield results opposite to the true relationship between $X_{1}^{(j)}$ and $X_{2}^{(j)}$ in each environment (cf. Figure 2 and Section C.1 for details).
We designate the first dimension, $X_{1}$, as the input feature, while the second dimension, $Y=X_{2}$, serves as the outcome variable. Importantly, the dataset meets our necessary criteria: We cannot infer the origin environment from a single input alone, as indicated by the overlap between the marginal distributions of $X_{1}$, obtained by projecting the entire sample onto the $x$-axis in Figure 2. Thus, this setting aligns with Criterion 2.2, and, additionally, corresponds to Criterion 2.3, since learning about the environment location $\boldsymbol{\mu}^{(j)}$ should improve prediction.
Figure 3: Experiment 1. Relative improvement of set-encoder (shown in I) approach versus baseline model (0 means no improvement is achieved) on toy example. We also show I (OOD) on OOD data. II depicts the relative improvement of the environment-oracle model compared to the baseline model. III demonstrates the relative improvement in predicting the environment when using contextual information compared to the absence of it. Sampling variation arises from using different seeds to partition the ID data into training, test and validation set.
@@ Results
As a first check of Criterion 2.2, we evaluate whether a set input provides additional information about the environment compared to a singleton input. This evaluation involves studying the classification accuracy in distinguishing between the environments. In Figure 2 we observe that the additional set input improves the ability to predict the environment significantly and the more samples we include, the better the prediction. As expected, a decrease in the distance between environments necessitates more samples to differentiate between environments. Interestingly, the particular choice of architecture for the permutation-invariant network does not seem to play a significant role for predicting the environment label well, as demonstrated in Appendix G.
Next, we assess the predictive capabilities of our approach across all possible scenarios of “leave-one-environment-out”. This involves training on all environments except one and treating the excluded environment as a novel OOD scenario. Here, we consider linear models to ensure an optimal inductive bias for the problem (non-linear models achieve similar results, as shown in Section C.3). We can see that Criterion 2.1, Criterion 2.2 and Criterion 2.3 are satisfied in Figure 3. Providing contextual information in the form of a set input increases the performance significantly compared to a baseline model in the ID as well as in the OOD setting (see I and I (OOD) in Figure 3). We also observe a slightly higher relative improvement when the environment label is directly provided (see II) compared to using the output of the set-encoder (see I). This aligns with our expectations, as the set input does not offer more information about the target value than the environment label itself.
Finally, we visualize the predictions of the baseline approach and our set-encoder approach in Figure 4. Our model captures and utilizes the characteristics of each environment for prediction. In contrast, the baseline approach struggles to discern between environments due to the significant overlap between environments, resulting in an inability to deal with environmental differences. Note that we obtained the best results on this problem by considering a class of linear models that aligns with the data-generating process. However, we observe that extrapolation performance drops when the considered models are overly complex and lack a strong inductive bias (see Section C.3).
4.3 Experiment 2: ProDAS Example
Setup
We utilize the ProDAS library [40] to generate high-dimensional image data that meets our dataset requirements. The dataset comprises objects of shape square and circle, exhibiting variations in their texture, background color, rotation, and size. Additionally, the background varies in color and texture, resulting in a complex scenario. We consider the task of predicting the object size. Difficulties arise due to the presence of distinct environments with varying characteristics. Specifically, depending on the environment, a constant is added to the observed object size to get the actual target variable that we aim to predict:
$Y_{\text{gt}}=Y_{\text{observed}}+j\cdot\text{const}_{1}$ (7)
Here, $j\in\{1,2,3,4\}$ denotes the environment, while $Y_{\text{gt}}$ represents the ground truth (or factual) size, obtained as a sum of the observed size $Y_{\text{observed}}$ (relative to the image frame) and a constant depending on $j$. The background color follows a normal distribution $\mathcal{N}(\boldsymbol{\mu}_{j};\boldsymbol{\Sigma})$ where the mean depends on the environment in the following way: $\boldsymbol{\mu}_{j}=\boldsymbol{\mu}_{0}+j\cdot\text{const}_{2}$. Here we assign a small value to $\text{const}_{2}$ to enforce the background distributions to overlap between different environments. Specifically, this construction implies that the relation between input $\mathbf{X}$ and target $Y$ differs across environments. This corresponds to Criterion 2.3. Notably, inferring the originating environment from a single sample is unattainable due to overlapping background distributions (corresponding to Criterion 2.2). Samples of different environments are shown in Appendix D. This example could be inspired by microscopy data where different microscopes correspond to distinct environments, each exhibiting its own characteristics. During training, we assume to have access to the ground truth value $Y_{\text{gt}}$.
Results
@@
In line with the results from the previous toy example, we can demonstrate a strong relative improvement in the ProDAS dataset, as depicted in Figure 5. All formal criteria are satisfied and a very significant improvement is achieved, both in the ID and the OOD setting, by considering the contextual information from the environment. Details for this experiment can be found in Appendix D.
4.4 Experiment 3: Colored MNIST
Setup
The ColoredMNIST dataset [4] is an extension of the standard MNIST dataset, wherein the number of classes is reduced to two classes (all standard labels $<5$ are assigned to new label 0, and all labels $\geq 5$ are the new label 1). Furthermore, there exists label noise, so only in 75% of all cases, the label can correctly be predicted from the input image. To make things more challenging, the image background can take two colors that are also associated with the image label. In the first environment, the association is 90% and in the second one 80%. Therefore, a baseline model would tend to utilize the background for prediction instead of the actual shape. However, in a third environment, the associations are reversed, so that a model based on the background color would achieve only 10% accuracy – worse than random.
This dataset implies a trade-off between predictive performance in ID domains versus robustness in OOD domains, as discussed in [42]. For instance, an invariant model that relies solely on an object’s shape would be robust to domain shift at the cost of diminished accuracy in the first two environments (75% instead of 80% or 90%). In contrast, a baseline model would achieve greater accuracy in the first domains (80% and 90%), but would fail dramatically in the third domain (only 10%).
@@ 4.4 Experiment 3: Colored MNIST
[…]
In-Distribution Out-of-Distribution
Baseline $\mathbf{84.6}\pm 0.3$ $10.2\pm 0.3$
Invariant $72.8\pm 0.9$ $\mathbf{73.1}\pm 0.2$
Selection (Ours) $\mathbf{84.1}\pm 0.3$ $\mathbf{73.1}\pm 0.2$
Selection (Baseline) $\mathbf{84.0}\pm 0.3$ $14.0\pm 0.4$
Bayes Optimal Classifier $\mathbf{85.0}$ $\mathbf{75.0}$
Table 2: Experiment 3. Mean and standard deviation of accuracy percentages across model types and domain settings over 5 runs. The features extracted by our model allow for improved OOD detection compared to the features of the baseline model. Thus, our model can perform a favorable selection between the baseline model in the ID setting and the invariant model in the OOD setting.
Results
Here, we assume the invariant model to be given (see Appendix E for details), but it could also be obtained by invariant learning, e.g. Invariant Risk Minimization [4]. With our novel environment detection approach (see Section 2.7) we can get the best of both worlds, circumventing the inherent trade-off. When identifying the ID setting, we utilize the baseline model that achieves the highest predictiveness within the observed environments. In case we detect the OOD setting, we employ the invariant model. We compare this kind of model selection due to the features $h_{\boldsymbol{\psi}}(\mathbf{S}^{(n)})$ inherent to our model versus the features extracted by the baseline model. The results can be found in Table 2. By utilizing the model selection based on the set-summary $h_{\boldsymbol{\psi}}(\mathbf{S}^{(n)})$, we nearly recover the ID accuracy while maintaining identical performance to the invariant model on OOD data. Evidently, the novel environment detection only works with set summaries. A feature extracted from a single sample does not provide enough information to reliably detect distribution shifts, leading to difficulties in effectively selecting between baseline and invariant model, as demonstrated in Table 2.
4.5 Experiment 4: Violated Criteria
Setup
We consider the PACS dataset [34], training our model on the Cartoons, Sketches, and Paintings environments, and assess its performance in the Art environment during testing. The dataset includes images with labels that we intend to predict. Regarding a second classification task, we delve into the OfficeHome dataset [55]. In line with the PACS dataset, we approach the classification problem, training across three specific environments, and subsequently evaluating a novel one as an out-of-distribution (OOD) scenario.
(a) Environment Art in PACS dataset. The environment is almost completely inferable from one input sample (Criterion 2.2 not satisfied). Conclusively our approach does not yield benefits.
Figure 6: Experiment 4. Examples where at least one of the necessary criteria is not satisfied and our approach cannot possibly yield benefits. For experimental details, see Appendix F.
Results
When the criteria are not met, no benefits can be achieved, even in the ID setting. This has been proven in Theorem 2.1 and we demonstrate it here for two scenarios empirically (see Figure 6). We find that Criterion 2.2 is not satisfied on the PACS dataset: As depicted in 5(b), the contextual environment model $f^{E\mid\mathbf{X},\mathbf{S}^{(n)}}$ does not perform better compared to the baseline environment model $f^{E\mid\mathbf{X}}$. Remarkably, a single example is sufficient to infer the source environment, allowing for a 99.7% accuracy in predicting the correct environment from an individual sample (see Appendix F). Since Criterion 2.2 is not fulfilled, we anticipate Criterion 2.3 also to be wrong. This is indeed the case as 5(b) depicts. Since the criteria are not met, we do not achieve any benefit over the baseline model, neither in the ID nor in the OOD setting, as demonstrated in 5(b).
On the OfficeHome dataset we find that Criterion 2.2 is not satisfied, while Criterion 2.3 is. Results are depicted in 5(b). We observe that the set input offers benefits for predicting the data originating environment corresponding to Criterion 2.3. However, even when providing the target classifier with the environment label (environment-oracle model), we do not achieve an improvement over the baseline model suggesting that Criterion 2.2 is not satisfied. As expected and depicted in 5(b), our method does not yield benefits compared to the baseline model.
4.6 Experiment 5: Failure Case Detection
Setup
Besides unfulfilled criteria, another reason why our approach might fail to reap benefits or degrade in performance is when the distribution shift requires extrapolation. This might be unattainable by the model. We demonstrate, using the BikeSharing dataset [17], that in cases where different seasons like summer or winter represent distinct environments, extrapolation might be necessary. For this dataset we consider the task of predicting the number of bikes rented across the day based on weather data. Details about the dataset and pre-processing steps can be found in Appendix H. We explore four scenarios, each entailing training on all seasons except one. We aim to assess the abilities of our model compared to a baseline model in detecting novel environments.
Results
In Table 3 we demonstrate that our approach is slightly superior compared to the baseline model in the ID settings. However, both the baseline and our approach experience performance degradation in the novel environments. To detect the novel environment (and consequentially potential failure cases), we compute the score as suggested in Section 2.7 and evaluate how well it distinguishes between ID versus OOD environment. We designate an independent ID test set and use the environment excluded during training (e.g., summer or winter) as the OOD set for evaluation. The area under the ROC-curve (AUROC) in Table 3 demonstrates that the score based on the summary vector provided by the permutation invariant network allows for a perfect novel environment detection whereas the standard approach fails in detecting the novel environment.
Table 3: Experiment 5. Performance comparison between our model and the baseline, broken down by target domain. We compare the performance in the ID and OOD setting (MSE), as well as their capability to detect a novel environment (AUROC). Both models fail in the OOD setting, but our model can detect with strong certainty when this is the case. We present the mean and standard deviation derived from 5 runs using different seeds for partitioning into training, validation, and test sets.
5 Conclusions
In this work, we introduce a novel approach enabling the extraction and utilization of contextual information from set inputs within a Domain Generalization (DG) setting. The set inputs originate from the same environment as the sample for which a prediction is required and are summarized into a vector via a permutation-invariant network. We formulate criteria that are necessary for our approach to yield benefits and are easy to verify. Empirically, we demonstrate that we can verify these criteria, which enables us to identify cases where our approach does not yield benefits. We showcase the merits of our approach on several datasets. Additionally, we demonstrate that novel environments can be detected, allowing for the identification of potential failure cases.
In our current framing, we have not yet employed regularization techniques, such as enforcing the set vector to contain all information about the environment in our composite model in Figure 1. In general, investigating inductive biases for the composite network architecture is an interesting future research avenue. Finally, even though we focused our experiments in standard supervised learning settings, our method could also be employed in other realms, such as domain disentanglement, data generation, or in combination with other DG methods. Thus, extending our method to different settings represents another exciting research direction.
Acknowledgments
Appendix A Theory
A.1 Insufficiency of Criteria 2 and 3 for Criteria 1
@@ A.2 Proof of Theorem
$0\leq I(\mathbf{Y};\mathbf{S}^{(n)}\mid\mathbf{X})=I(\mathbf{S}^{(n)};\mathbf{Y},\mathbf{X})-I(\mathbf{X},\mathbf{S}^{(n)})\leq 0$ (16)
@@
Section A.2 can be extended to the conditional mutual information if $A\perp B\mid D$ and $A\perp B\mid D,C$:
@@
Appendix B Experiments: General Remarks
@@
$\mathcal{R}_{\text{II}}=\frac{\mathcal{L}(f^{E\mid\mathbf{X},\mathbf{S}^{(n)}})-\mathcal{L}(f^{E\mid\mathbf{X}})}{\mathcal{L}(f^{E\mid\mathbf{X},\mathbf{S}^{(n)}})}$ (26)
@@
$\mathcal{R}_{\text{III}}=\frac{\mathcal{L}(f^{\mathbf{Y}\mid\mathbf{X},E})-\mathcal{L}(f^{\mathbf{Y}\mid\mathbf{X}})}{\mathcal{L}(f^{\mathbf{Y}\mid\mathbf{X},E})}$ (27)
@@
Due to the large amount of settings, we did only little hyper-parameter optimization (we looked into batch size, learning rate, and network size). For a given dataset we optimized only on one scenario where an environment is left out during training. The found hyper-parameters were then applied on all other scenarios. To ensure that the baseline model is comparable to ours, we ensure that the inference network (and feature extractor) in Figure 1 have a comparable number of parameters as the baseline model. In all cases, the set-encoder is kept simple and its hyper-parameters are selected for optimal performance of the contextual environment predictor $f^{E\mid\mathbf{X},\mathbf{S}^{(n)}}$. For an overview, see Table 5. Throughout all experiments, we employ a mean-pooling operation.
@@
Appendix C Experiment 1: Details
C.1 Data Generation
Simpson’s Paradox [45, 57] describes a statistical phenomenon wherein several groups of data exhibit a trend, which reverses when the groups are combined. There are several famous real-world examples of Simpson’s Paradox, such as a study examining a gender bias in the admission process of UC Berkeley [6] or an evaluation of the efficacy of different treatments for kidney stones [9].
In order to replicate this, we create a dataset as a mixture of 2D multivariate normal distributions, with the intent of using the first dimension as a feature, and the second as a regression target. Unless otherwise specified, we generate the data by taking an equal number of samples from each mixture component, defining the environment as a one-hot vector over the mixture components.
The mixture components are chosen to lie on a trend line that is opposite to the trend within each mixture. We achieve this by using a negative global trend, and choosing the covariance matrix of each mixture as a scaled and rotated identity matrix with opposite trend.

### 21 Feb 2024 v2

@@ A.1 Model Overview
Appendix B Additional Experiment: ProDAS
@@
We utilize the ProDAS library [Müller et al., 2023] to generate high-dimensional image data that meets our dataset requirements. The dataset comprises objects of shape square and circle, exhibiting variations in their texture, background color, rotation, and size. Additionally, the background varies in color and texture, resulting in a complex scenario. For examples see Figure 9. We consider the task of predicting the object size. Difficulties arise due to the presence of distinct environments with varying characteristics. Specifically, depending on the environment, a constant is added to the observed object size to get the actual target variable that we aim to predict:
$Y_{\text{gt}}=Y_{\text{observed}}+j\cdot\text{const}_{1}$ (3)
Here, $j\in\{1,2,3,4\}$ denotes the environment, while $Y_{\text{gt}}$ represents the ground truth (or factual) size, obtained as a sum of the observed size $Y_{\text{observed}}$ (relative to the image frame) and a constant depending on $j$. The background color follows a normal distribution $\mathcal{N}(\boldsymbol{\mu}_{j};\boldsymbol{\Sigma})$ where the mean depends on the environment in the following way: $\boldsymbol{\mu}_{j}=\boldsymbol{\mu}_{0}+j\cdot\text{const}_{2}$. Here we assign a small value to $\text{const}_{2}$ to enforce the background distributions to overlap between different environments. Specifically, this construction implies that the relation between input $\mathbf{X}$ and target $Y$ differs across environments. This corresponds to Criterion 2.3. Notably, inferring the originating environment from a single sample is unattainable due to overlapping background distributions (corresponding to Criterion 2.2). Samples of different environments are shown in Appendix F. This example could be inspired by microscopy data where different microscopes correspond to distinct environments, each exhibiting its own characteristics. During training, we assume to have access to the ground truth value $Y_{\text{gt}}$.
@@ Results
In line with the results from the previous toy example, we can demonstrate a strong relative improvement in the ProDAS dataset, as depicted in Figure 6. All formal criteria are satisfied and a very significant improvement is achieved, both in the ID and the OOD setting, by considering the contextual information from the environment. Additional details for this experiment can be found in Appendix F.
@@
Appendix C Theory
C.1 Generalization of Theorem 2.1 to Noisy Environments
Theorem C.1.
In addition to Theorem 2.1, the following holds:
  • (d)
The proof can be found in subsection C.3.
C.2 Insufficiency of Criteria 2 and 3 for Criterion 1
@@
$0\leq I(\mathbf{Y};\mathbf{S}^{(n)}\mid\mathbf{X})=I(\mathbf{S}^{(n)};\mathbf{Y},\mathbf{X})-I(\mathbf{X};\mathbf{S}^{(n)})\leq 0$ (12)
@@
subsection C.3 can be extended to the conditional mutual information if $A\perp B\mid D$ and $A\perp B\mid D,C$:
@@
Appendix D Experiments: General Remarks
@@
$\mathcal{R}_{\text{II}}=\frac{\mathcal{L}(f^{E\mid\mathbf{X},\mathbf{S}^{(n)}})-\mathcal{L}(f^{E\mid\mathbf{X}})}{\mathcal{L}(f^{E\mid\mathbf{X}})}$ (22)
@@
$\mathcal{R}_{\text{III}}=\frac{\mathcal{L}(f^{\mathbf{Y}\mid\mathbf{X},E})-\mathcal{L}(f^{\mathbf{Y}\mid\mathbf{X}})}{\mathcal{L}(f^{\mathbf{Y}\mid\mathbf{X}})}$ (23)
@@
Due to the large amount of settings, we did only little hyper-parameter optimization (we looked into batch size, learning rate, and network size). For a given dataset we optimized only on one scenario where an environment is left out during training. The found hyper-parameters were then applied to all other scenarios. To ensure that the baseline model is comparable to ours, we ensure that the inference network (and feature extractor) in Figure 1 have a comparable number of parameters as the baseline model. In all cases, the set-encoder is kept simple and its hyper-parameters are selected for optimal performance of the contextual environment predictor $f^{E\mid\mathbf{X},\mathbf{S}^{(n)}}$. For an overview, see Table 6. Throughout all experiments, we employ a mean-pooling operation.
@@
Appendix E Experiment 1: Details
E.1 Data Generation
Simpson’s Paradox [Peters et al., 2017, von Kügelgen et al., 2021] describes a statistical phenomenon wherein several groups of data exhibit a trend, which reverses when the groups are combined. There are several famous real-world examples of Simpson’s Paradox, such as a study examining a gender bias in the admission process of UC Berkeley [Bickel et al., 1975] or an evaluation of the efficacy of different treatments for kidney stones [Charig et al., 1986].
To replicate this, we create a dataset inspired by an illustration of Simpson’s Paradox on Wikipedia sim . The dataset consists of a mixture of 2D multivariate normal distributions, with the intent of using the first dimension as a feature, and the second as a regression target. Unless otherwise specified, we generate the data by taking an equal number of samples from each mixture component, defining the environment as a one-hot vector over the mixture components.
The mixture components are chosen to lie on a trend line that is opposite to the trend within each mixture. We achieve this by using a negative global trend and choosing the covariance matrix of each mixture as a scaled and rotated identity matrix with an opposite trend.